What is the detail solution of this question:
In a class of 50 students the boys obtained an average grade of 80 points for the Maths
test and the girls obtained an average grade of 90 points for the same test. If the boys had scored 5%
more on average and the girls had scored 10% more on average, then the whole class would have
got 350 extra points. How many boys are there in this class?

Answer by stanbon(75887)

In a class of 50 students the boys obtained an average grade of 80 points for the Maths test and the girls obtained an average grade of 90 points for the same test.
-----
# of boys + # of girls = 50
b + g = 50
-------------------------------
Points DATA:
Boy's point total:: 80b
Girl's point total: 90g
Original total points:: 80b + 90g
------------------------------------
If the boys had scored 5% more on average and the girls had scored 10% more on average, then the whole class would have got 350 extra points.
Boys point total would be 1.05*80b = 84b
Girls point total would be 1.10*90g = 99g
New number of points = 84b+99g
----

How many boys are there in this class?
New number of points - original # of points = 350
[84b + 99g] - [80b+90g] = 350
----
Two equations from the DATA::
4b + 9g = 350
b + g = 50
---------------------
Modify for elimination:
4b + 9g = 350
9b + 9g = 450
---
5b = 100
b = 20 (# of boys)
===========================
Cheers,
Stan H.